OpenStudy (anonymous):

For the following function: f(x)=x^4-3 Find the maximum number of real zeros, and Use Descartes's rule of signs to determine how many positive and how many negative zeros the function has. You do not need to find the zeros.

OpenStudy (anonymous):

1 negative, 1 positive, 2 complex conjugates

OpenStudy (anonymous):

f(x)=x^4-3 has 1 sign change, 1 positive real zero f(-x)=(-x)^4-3=x^4-3, 1 sign change, 1 negative real zero the other 2 are complex conjuages